OpenStudy (anonymous):

Simplify. Express the product as a radical expression. (x^4)^3/5

OpenStudy (anonymous):

HELP PLEASE?!

OpenStudy (anonymous):

do u know the rule for power of a power: \(\large (b^a)^c=b^{ac} \) ??? apply that then use the fact that: \(\large b^{\frac{p}{q}}=\sqrt[q]{b^p} \)

OpenStudy (anonymous):

no...